please show all work
Find the indicated partial derivatives. F(x, y, z) = Py(6, 1, -1) X + y + (6, 1, -1) Submit Answer
Answer #1


5 SE T = g z(H+9) (17) +9 fy (6, 1, 1) = z (8tht) {tk = they zrethtk) h Ethic = zlethal (7) - (D&th) SPAI AR h =ff atytz) - (